hollidayli:

What are the solutions to log3x + log3(x2 + 2) = 1 + 2log3x? x = –2 x = –1 x = 1 x = 2 There is no true solution.

jhonyy9:

use the properties of logs log a + log b = log a×b

surjithayer:

\[\log_{3} x+\log_{3} (x^2+2)=1+2\log_{3} x\] \[\log_{3} x-2\log_{3} x+\log_{3} (x^2+2)=1\] \[-\log_{3} x+\log_{3} (x^2+2)=\log_{3} 3\] \[\log_{3} \frac{ x^2+2 }{ x }=\log_{3} 3\] \[\frac{ x^2+2 }{ x }=3\] \[x^2+2=3x\] \[x^2-3x+2=0\] ?